# Equivalence Relations on Vertex Operator Algebras, II: Witt Equivalence and Orbifolds

### v1

@@ 4.1. Orbifold Equivalence
As an example, consider the vertex operator algebra $V=\smash{\mathsf{A}_{1,1}^{G(E_{8})}}$ described in 4.6, and let $\mathcal{T}$ be the diagonal RCFT constructed from $V$. It is believed that any finite-rank fusion category symmetry acting on $\mathcal{T}$ can be embedded into its (finite-rank) category of Verlinde lines, $\operatorname{Ver}(\mathcal{T}/V,V)$. Indeed, suppose there is some fusion category $\mathcal{F}$ that does not commute with the chiral algebra $V$. Then one can form the subalgebra $V^{\mathcal{F}}$ of operators in $V$ that commute with the lines in $\mathcal{F}$. Such a subalgebra $V^{\mathcal{F}}$ would ostensibly define a new positive, strongly rational, $c=1$ vertex operator algebra, contradicting the conjectured classification in 3.26.
@@ 4.2. Interface Equivalence
Definitionph 4.9 (Interface Equivalence).
@@
We can apply this same way of thinking to two-dimensional interfaces, as opposed to 2d QFTs. Indeed, a topological interface $\mathcal{I}$ between the 3d TQFTs $\operatorname{Rep}(V)$ and $\operatorname{Rep}(W)$ always supports a fusion category $\mathcal{F}$ of topological line operators on its world-volume, and it is always possible to reach any other topological interface $\mathcal{I}^{\prime}$ by performing a 1-gauging of $\mathcal{F}$ [RSS23]. Using the same kinds of arguments as in the previous paragraph, the existence of the line interface $I$ between the $V$ and $W$ boundaries on which $\mathcal{I}$ terminates guarantees, for any other topological interface $\mathcal{I}^{\prime}$, the existence of a line interface $I^{\prime}$ between the $V$ and $W$ boundaries on which $\mathcal{I}^{\prime}$ can terminate. ∎
Propositionph 4.11.
@@
$\xrightarrow{\mathrm{fuse}}$$\operatorname{Rep}(V^{\prime})$$\operatorname{Rep}(V)$$V$$V^{\prime}$
Figure 13. Depiction of the proof of 4.11.
Remarkph 4.12.
It is tempting to assert, in the spirit of [DLWW24], that the converse of 4.11 holds as well. While this certainly seems conceivable to us, it is far from obvious. We will offer further intuition below, in the more familiar (to physicists) context of full RCFTs.
@@
Again, by using the trick of “performing an orbifold in half of spacetime”, we can derive the following analog of 4.11.
Propositionph 4.14.
@@
Remarkph 4.15.
One might come away from [DLWW24] with the impression that the converse to 4.14 holds. Indeed, op. cit. shows that if $I$ is a line interface between two RCFTs $\mathcal{T}$ and $\mathcal{T}^{\prime}$, then the fusion of $I$ with its orientation reversal, $A\coloneqq I^{\ast}\otimes I$, defines a (non-simple) topological line operator inside $\mathcal{T}$, and in fact a finite quantum dimension algebra object that can be gauged to reach $\mathcal{T}^{\prime}$. However, in this argument, no reference is made to the ambient category in which $A$ sits. Even though $A$ decomposes into finitely many simple line operators by virtue of its finite quantum dimension, it is not clear that the category generated by the simple lines appearing in $A$ has finite rank.
Indeed, returning to the example discussed at the end of 4.8, if one composes the half-space gauging line interface between $\mathcal{T}$ and $\mathrm{SU}(2)_{1}$ with the half-space gauging line interface between $\mathrm{SU}(2)_{1}$ and $\smash{\mathcal{T}^{\mathrm{circ}}_{\sqrt{2}p/q}}$, then one obtains an interface $I$ between $\mathcal{T}$ and $\smash{\mathcal{T}^{\mathrm{circ}}_{\sqrt{2}p/q}}$. Even though $I^{\ast}\otimes I$ must contain finitely many simple topological lines, evidently the category these simple lines generate must have infinite rank for most choices of $p$ and $q$.
@@ 4.3. Inner Orbifold Equivalence
Definitionph 4.19 (Inner Orbifold Equivalence for RCFTs).
@@
We have chosen the label “inner” because, in some special cases, inner orbifold equivalence reduces to a (group-like) orbifold construction under inner automorphisms.
Finally, it follows directly from the definition that inner orbifold equivalence (like orbifold equivalence, see 4.4) reduces to rational equivalence of lattices if we consider lattice vertex operator algebras. Of course, this is by design.
@@ 5.1. Necessary Condition
By 4.16, $V$ is orbifold equivalent to $V^{\prime}$ and $W$ is orbifold equivalent to $W^{\prime}$. Then, by 5.1, $V$ and $V^{\prime}$ are Witt equivalent as strongly rational vertex operator algebras, and so are $W$ and $W^{\prime}$. This is the definition of ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ being Witt equivalent as RCFTs. ∎
@@
By 4.17, $V$ is interface equivalent to $V^{\prime}$ and $W$ is interface equivalent to $W^{\prime}$. By 5.2, $V$ is Witt equivalent to $V^{\prime}$ and $W$ is Witt equivalent to $W^{\prime}$, which is by definition what it means for ${{}_{V}}\mathcal{H}_{W}$ to be Witt equivalent to ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$. ∎
Thus, we see that Witt inequivalence is a kind of obstruction to being orbifold or interface equivalent. In particular, by 3.14, one has a criterion for testing whether two RCFTs may be related by a sequence of generalized orbifolds, or separated by a topological line interface: one computes the triple $(c_{L},c_{R},[\operatorname{Rep}(V)])=(c_{L},c_{R},[\operatorname{Rep}(W)])$ corresponding to ${{}_{V}}\mathcal{H}_{W}$ and the triple $(c_{L}^{\prime},c_{R}^{\prime},[\operatorname{Rep}(V^{\prime})])=(c_{L}^{\prime},c_{R}^{\prime},[\operatorname{Rep}(W^{\prime})])$ corresponding to ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ and simply checks whether or not they are equal. (Here, e.g., $[\operatorname{Rep}(V)]$ is the Witt class of the modular tensor category $\operatorname{Rep}(V)$.) If the triples are equal, then ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ might be related by an orbifold; if they are not equal, then the theories are definitively not related by an orbifold.
@@ 5.2. Converse Conjectures
Recall from 4.4 that orbifold equivalence generalizes rational equivalence of even lattices. Moreover, in the context of lattices, in addition to rational equivalence implying Witt equivalence, also the converse holds (see 2.6). Together with 3.1, we obtain:
@@
We shall not establish these conjectures in this work. This is probably a very difficult problem, which, as we shall see in 5.16 below, is likely at least as difficult as proving the moonshine uniqueness, 2.20.
@@
On the other hand, assume 5.11 and consider two positive, strongly rational vertex operator algebras $V$ and $V^{\prime}$ belonging to the same Witt class. By 2.13, there is a unique anisotropic representative $\mathcal{C}_{\mathrm{an}}$ of the Witt class $[\operatorname{Rep}(V)]=[\operatorname{Rep}(V^{\prime})]$. Now, by 2.18, $V$ and $V^{\prime}$ can be conformally embedded into vertex operator algebras $W$ and $W^{\prime}$, respectively, both belonging to the bulk genus $(\mathcal{C}_{\mathrm{an}},c(V))=(\mathcal{C}_{\mathrm{an}},c(V^{\prime}))$. 5.11 then implies that $W$ is orbifold equivalent to $W^{\prime}$, and because $V$ is orbifold equivalent to $W$ (and $V^{\prime}$ orbifold equivalent to $W^{\prime}$), it follows that $V$ is orbifold equivalent to $V^{\prime}$, and hence that 5.6 is true. ∎
@@
For $c=0,8,16$ the bulk genus $([\operatorname{Vect}],c)$ consists only of lattice vertex operator algebras. These are then all (inner) orbifold equivalent because, e.g., the $2$-neighborhood graph of the corresponding positive-definite, even, unimodular lattices is connected and lattice neighborhood can clearly be lifted to neighborhood of lattice vertex operator algebras. See 5.21 below for more details.
For $c=24$, the situation is more difficult because most vertex operator algebras in the bulk genus $([\operatorname{Vect}],24)$ are not lattice vertex operator algebras. Nonetheless, the main result of [ELMS21, MS23] establishes that, assuming 2.20 on the uniqueness of the moonshine module, all strongly rational, holomorphic vertex operator algebras of central charge $24$ can be obtained as cyclic orbifold constructions of the Leech lattice vertex operator algebra $V_{\Lambda}$ associated with generalized deep holes and are thus all orbifold equivalent to one another. (See also [HM23] for another approach with the same consequence.)
@@
Assuming 2.20 on the uniqueness of the moonshine module, 5.6 is true for the infinitely many positive, strongly rational vertex operator algebras belonging to the Witt classes $([\operatorname{Vect}],c)$ with $c\leq 24$.
@@
Assuming 2.20 on the uniqueness of the moonshine module, 5.8 is true for the infinitely many positive RCFTs belonging to the Witt classes $(c_{L},c_{R},[\operatorname{Vect}])$ with $c_{L},c_{R}\leq 24$.
@@
Indeed, it is not difficult to show that if such a “fake” moonshine module $V$ existed, it could not be orbifold equivalent under iterated cyclic orbifold constructions to any of the $71$ Schellekens vertex operator algebras. (And likely, this argument can be extended to the more general notion of orbifold equivalence given in 4.3.)
@@ 5.3. Partial Converse
Establishing 5.6 (or 5.11) in full generality appears to be a very difficult problem. For example, as we just saw in 5.16, as a very narrow special case, one would likely obtain the uniqueness of the moonshine module, 2.20. Instead, we prove the following partial result, which shows that at least strongly Witt equivalent vertex operator algebras are necessarily orbifold equivalent, and in fact inner orbifold equivalent.
@@
Suppose $V$ and $V^{\prime}$ are inner orbifold equivalent, recalling the notation from 4.18. In each intermediate step, since $W_{i}$ and $V_{i}$ have the same Heisenberg commutant $C$, we can choose a Cartan subalgebra of $W_{i}$ that is also a Cartan subalgebra of $V_{i}$. Indeed, any Cartan subalgebra of $W_{i}$ can be extended to a Cartan subalgebra of $V_{i}$, but the condition on the Heisenberg commutants implies that these Cartan subalgebras have the same dimension.
@@
The proof also shows that while in general two orbifold equivalent vertex operator algebras cannot be connected “in one step” (see 4.6), this is the case for inner orbifold equivalence (cf. 2.5).
@@
Note that even inner orbifold equivalence is in general weaker than hyperbolic equivalence.
@@ 5.4. Examples
Let $\mathcal{C}_{\mathrm{an}}$ be an anisotropic unitary modular tensor category with $\operatorname{rk}(\mathcal{C}_{\mathrm{an}})\leq 4$ [RSW09] other than $\smash{\overline{(A_{1},5)}_{\frac{1}{2}}}$ (see op. cit. for the definition of the latter). Further, let $c\in\mathbb{Q}$ be the unique rational number satisfying $0<c\leq 8$ and $e^{2\pi\mathrm{i}c/8}=e^{2\pi\mathrm{i}c(\mathcal{C}_{\mathrm{an}})/8}$. Here, we recall from [MR24, LABEL:MR1eqn:chiralcentralcharge] the definition of the chiral central charge $c(\mathcal{C}_{\mathrm{an}})\in\mathbb{Q}/8\mathbb{Z}$ of $\mathcal{C}_{\mathrm{an}}$.
@@
Using 4.16, this result can be extended to full RCFTs. In particular, one finds that the entire Witt class $(c_{L},c_{R},[\mathcal{C}_{\mathrm{an}}])=(c,c,[\mathcal{C}_{\mathrm{an}}])$ of the canonical diagonal RCFT built on top of $V$ is connected by orbifolds. For example, every RCFT that is Witt equivalent to the $G_{2,1}$ Wess–Zumino–Witten model can be reached by performing a sequence of orbifolds on the latter.

### v2

@@ 4.1. Orbifold Equivalence
Remark 4.4.
From [Rie22], it is known that $W_{i}$ can always be obtained as the fixed-point subalgebra with respect to a hypergroup action on $V_{i}$. Thus, one can think of 4.3 as a generalization of 4.2, which captures what it means for two strongly rational vertex operator algebras to be related by hypergroup orbifolds (as opposed to orbifolds by abelian finite groups).
@@
As an example, consider the vertex operator algebra $V=\smash{\mathsf{A}_{1,1}^{G(E_{8})}}$ described in 4.7, and let $\mathcal{T}$ be the diagonal RCFT constructed from $V$. It is believed that any finite-rank fusion category symmetry acting on $\mathcal{T}$ can be embedded into its (finite-rank) category of Verlinde lines, $\operatorname{Ver}(\mathcal{T}/V,V)$. Indeed, suppose there is some fusion category $\mathcal{F}$ that does not commute with the chiral algebra $V$. Then one can form the subalgebra $V^{\mathcal{F}}$ of operators in $V$ that commute with the lines in $\mathcal{F}$. Such a subalgebra $V^{\mathcal{F}}$ would ostensibly define a new positive, strongly rational, $c=1$ vertex operator algebra, contradicting the conjectured classification in 3.26.
@@ 4.2. Interface Equivalence
Definitionph 4.10 (Interface Equivalence).
@@
We can apply this same way of thinking to $2$-dimensional interfaces, as opposed to 2d QFTs. Indeed, a topological interface $\mathcal{I}$ between the 3d TQFTs $\operatorname{Rep}(V)$ and $\operatorname{Rep}(W)$ always supports a fusion category $\mathcal{F}$ of topological line operators on its world-volume, and it is always possible to reach any other topological interface $\mathcal{I}^{\prime}$ by performing a 1-gauging of $\mathcal{F}$ [RSS23]. Using the same kinds of arguments as in the previous paragraph, the existence of the line interface $I$ between the $V$ and $W$ boundaries on which $\mathcal{I}$ terminates guarantees, for any other topological interface $\mathcal{I}^{\prime}$, the existence of a line interface $I^{\prime}$ between the $V$ and $W$ boundaries on which $\mathcal{I}^{\prime}$ can terminate. ∎
Propositionph 4.12.
@@
[…]
Figure 14. Depiction of the proof of 4.12.
Remarkph 4.13.
It is tempting to assert, in the spirit of [DLWW24], that the converse of 4.12 holds as well. While this certainly seems conceivable to us, it is far from obvious. We will offer further intuition below, in the more familiar (to physicists) context of full RCFTs.
@@
Again, by using the trick of “performing an orbifold in half of spacetime”, we can derive the following analog of 4.12.
Propositionph 4.15.
@@
Remarkph 4.16.
One might come away from [DLWW24] with the impression that the converse to 4.15 holds. Indeed, op. cit. shows that if $I$ is a line interface between two RCFTs $\mathcal{T}$ and $\mathcal{T}^{\prime}$, then the fusion of $I$ with its orientation reversal, $A\coloneqq I^{\ast}\otimes I$, defines a (non-simple) topological line operator inside $\mathcal{T}$, and in fact a finite quantum dimension algebra object that can be gauged to reach $\mathcal{T}^{\prime}$. However, in this argument, no reference is made to the ambient category in which $A$ sits. Even though $A$ decomposes into finitely many simple line operators by virtue of its finite quantum dimension, it is not clear that the category generated by the simple lines appearing in $A$ has finite rank.
Indeed, returning to the example discussed at the end of 4.9, if one composes the half-space gauging line interface between $\mathcal{T}$ and $\mathrm{SU}(2)_{1}$ with the half-space gauging line interface between $\mathrm{SU}(2)_{1}$ and $\smash{\mathcal{T}^{\mathrm{circ}}_{\sqrt{2}p/q}}$, then one obtains an interface $I$ between $\mathcal{T}$ and $\smash{\mathcal{T}^{\mathrm{circ}}_{\sqrt{2}p/q}}$. Even though $I^{\ast}\otimes I$ must contain finitely many simple topological lines, evidently the category these simple lines generate must have infinite rank for most choices of $p$ and $q$.
@@ 4.3. Inner Orbifold Equivalence
Definitionph 4.20 (Inner Orbifold Equivalence).
@@
We have chosen the label “inner” because, in some special cases, inner orbifold equivalence reduces to a (group-like) orbifold construction under inner automorphisms (see 8.6 and 8.7).
Finally, it follows directly from the definition that inner orbifold equivalence (like orbifold equivalence, see 4.5) reduces to rational equivalence of lattices if we consider lattice vertex operator algebras. Of course, this is by design.
@@ 5.1. Necessary Condition
By 4.17, $V$ is orbifold equivalent to $V^{\prime}$ and $W$ is orbifold equivalent to $W^{\prime}$. Then, by 5.1, $V$ and $V^{\prime}$ are Witt equivalent as strongly rational vertex operator algebras, and so are $W$ and $W^{\prime}$. This is the definition of ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ being Witt equivalent as RCFTs. ∎
@@
By 4.18, $V$ is interface equivalent to $V^{\prime}$ and $W$ is interface equivalent to $W^{\prime}$. By 5.2, $V$ is Witt equivalent to $V^{\prime}$ and $W$ is Witt equivalent to $W^{\prime}$, which is by definition what it means for ${{}_{V}}\mathcal{H}_{W}$ to be Witt equivalent to ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$. ∎
Thus, we see that Witt inequivalence is a kind of obstruction to being orbifold or interface equivalent. In particular, by 3.14, one has a criterion for testing whether two RCFTs may be related by a sequence of generalized orbifolds, or separated by a topological line interface: one computes the triple $([\operatorname{Rep}(V)],c_{L},c_{R})=([\operatorname{Rep}(W)],c_{L},c_{R})$ corresponding to ${{}_{V}}\mathcal{H}_{W}$ and the triple $([\operatorname{Rep}(V^{\prime})],c_{L}^{\prime},c_{R}^{\prime})=([\operatorname{Rep}(W^{\prime})],c_{L}^{\prime},c_{R}^{\prime})$ corresponding to ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ and simply checks whether or not they are equal. (Here, e.g., $[\operatorname{Rep}(V)]$ is the Witt class of the modular tensor category $\operatorname{Rep}(V)$.) If the triples are equal, then ${{}_{V}}\mathcal{H}_{W}$ and ${{}_{V^{\prime}}}\mathcal{H}^{\prime}_{W^{\prime}}$ might be related by an orbifold; if they are not equal, then the theories are definitively not related by an orbifold.
@@ 5.2. Converse Conjectures
Recall from 4.5 that orbifold equivalence generalizes rational equivalence of even lattices. Moreover, in the context of lattices, in addition to rational equivalence implying Witt equivalence, also the converse holds (see 2.6). Together with 3.1, we obtain:
@@
We shall not establish these conjectures in this work. This is probably a very difficult problem, which, as we shall see in 5.16 below, is likely at least as difficult as proving the moonshine uniqueness, 2.21.
@@
On the other hand, assume 5.11 and consider two positive, strongly rational vertex operator algebras $V$ and $V^{\prime}$ belonging to the same Witt class. By 2.14, there is a unique anisotropic representative $\mathcal{C}_{\mathrm{an}}$ of the Witt class $[\operatorname{Rep}(V)]=[\operatorname{Rep}(V^{\prime})]$. Now, by 2.19, $V$ and $V^{\prime}$ can be conformally embedded into vertex operator algebras $W$ and $W^{\prime}$, respectively, both belonging to the bulk genus $(\mathcal{C}_{\mathrm{an}},c(V))=(\mathcal{C}_{\mathrm{an}},c(V^{\prime}))$. 5.11 then implies that $W$ is orbifold equivalent to $W^{\prime}$, and because $V$ is orbifold equivalent to $W$ (and $V^{\prime}$ orbifold equivalent to $W^{\prime}$), it follows that $V$ is orbifold equivalent to $V^{\prime}$, and hence that 5.6 is true. ∎
@@
For $c=0,8,16$ the bulk genus $(\operatorname{Vect},c)$ consists only of lattice vertex operator algebras. These are then all (inner) orbifold equivalent because, e.g., the $2$-neighborhood graph of the corresponding positive-definite, even, unimodular lattices is connected and lattice neighborhood can clearly be lifted to neighborhood of lattice vertex operator algebras. See 5.21 below for more details.
For $c=24$, the situation is more difficult because most vertex operator algebras in the bulk genus $(\operatorname{Vect},24)$ are not lattice vertex operator algebras. Nonetheless, the main result of [ELMS21, MS23] establishes that, assuming 2.21 on the uniqueness of the moonshine module, all strongly rational, holomorphic vertex operator algebras of central charge $24$ can be obtained as cyclic orbifold constructions of the Leech lattice vertex operator algebra $V_{\Lambda}$ associated with generalized deep holes and are thus all orbifold equivalent to one another. (See also [HM23] for another approach with the same consequence.)
@@
Assuming 2.21 on the uniqueness of the moonshine module, 5.6 is true for the infinitely many positive, strongly rational vertex operator algebras belonging to the Witt classes $([\operatorname{Vect}],c)$ with $c\leq 24$.
@@
Assuming 2.21 on the uniqueness of the moonshine module, 5.8 is true for the infinitely many positive RCFTs belonging to the Witt classes $([\operatorname{Vect}],c_{L},c_{R})$ with $c_{L},c_{R}\leq 24$.
@@
Indeed, it is not difficult to show that if such a “fake” moonshine module $V$ existed, it could not be orbifold equivalent under iterated cyclic orbifold constructions (see Section 8 below) to any of the $71$ Schellekens vertex operator algebras. (And likely, this argument can be extended to the more general notion of orbifold equivalence given in 4.3.)
@@ 5.3. Partial Converse
Establishing 5.6 (or 5.11) in full generality appears to be a very difficult problem. For example, as we just saw in 5.16, as a very narrow special case, one would likely obtain the uniqueness of the moonshine module, 2.21. Instead, we prove the following partial result, which shows that at least strongly Witt equivalent vertex operator algebras are necessarily orbifold equivalent, and in fact inner orbifold equivalent.
@@
Suppose $V$ and $V^{\prime}$ are inner orbifold equivalent, recalling the notation from 4.19. In each intermediate step, since $W_{i}$ and $V_{i}$ have the same Heisenberg commutant $C$, we can choose a Cartan subalgebra of $W_{i}$ that is also a Cartan subalgebra of $V_{i}$. Indeed, any Cartan subalgebra of $W_{i}$ can be extended to a Cartan subalgebra of $V_{i}$, but the condition on the Heisenberg commutants implies that these Cartan subalgebras have the same dimension.
@@
The proof also shows that while in general two orbifold equivalent vertex operator algebras cannot be connected “in one step” (see 4.7), this is the case for inner orbifold equivalence (cf. 2.5).
@@
Note that even inner orbifold equivalence is in general weaker than hyperbolic equivalence. In Section 8, in the case of holomorphic vertex operator algebras $V$ and $V^{\prime}$, we shall make precise what we have to additionally require in order to find a notion of orbifold equivalence (namely inner $p$-neighborhood) that is equivalent to hyperbolic equivalence, at least conjecturally.
@@ 5.4. Examples
In Section 8, we discuss in detail a special case of orbifold equivalence, namely $n$-neighborhood or being related by a $\mathbb{Z}_{n}$-orbifold construction. The above is an example of this. In Section 8.5, we consider the holomorphic vertex operator algebras of central charge $c=24$.
@@
Let $\mathcal{C}_{\mathrm{an}}$ be an anisotropic unitary modular tensor category with $\operatorname{rk}(\mathcal{C}_{\mathrm{an}})\leq 4$ [RSW09] other than $\smash{\overline{(A_{1},5)}_{1/2}}$ (see op. cit. for the definition of the latter). Further, let $c\in\mathbb{Q}$ be the unique rational number satisfying $0<c\leq 8$ and $e^{2\pi ic/8}=e^{2\pi ic(\mathcal{C}_{\mathrm{an}})/8}$. Here, we recall from [MR24, LABEL:MR1eqn:chiralcentralcharge] the definition of the chiral central charge $c(\mathcal{C}_{\mathrm{an}})\in\mathbb{Q}/8\mathbb{Z}$ of $\mathcal{C}_{\mathrm{an}}$.
@@
Using 4.17, this result can be extended to full RCFTs. In particular, one finds that the entire Witt class $([\mathcal{C}_{\mathrm{an}}],c_{L},c_{R})=([\mathcal{C}_{\mathrm{an}}],c,c)$ of the canonical diagonal RCFT built on top of $V$ is connected by orbifolds. For example, every RCFT that is Witt equivalent to the $G_{2,1}$ Wess–Zumino–Witten model can be reached by performing a sequence of orbifolds on the latter.